OpenStudy (anonymous):

absolute value |x - 8| - 17 = -16

OpenStudy (anonymous):

\[|x-8|-17=-16\\|x-8|=1\\x-8=\pm1\\x=8\pm1\\x\in\{7,9\}\]

OpenStudy (anonymous):

Look at @oldrin.bataku response.